OpenStudy (anonymous):

5 square root of 7 + 12 square root of 6 - 10 square root of 7 - 5 square root of 6

OpenStudy (anonymous):

\[\Large 5\sqrt 7 + 12\sqrt 6 - 10\sqrt 7 - 5\sqrt 6 = 7\sqrt 6 - 5\sqrt 7 \]